OpenStudy (loli):

One bus left Denver at 3 pm going west at 40 mph and another left at 5 pm going west at 60 mph. How many hours will it take the second bus to catch up to the first? Round to the nearest tenth if necessary.

OpenStudy (radar):

When they have traveled the same distance, they will be at the same place. At 5 PM, bus1 has already traveled 80 miles so starting at d(bus1)=80 +40t bus2 d=60t They meet when the distances are equal. 80 + 40t = 60t 20t=80 t=4 hrs At 9:PM they meet (5PM + 4)

OpenStudy (radar):

I may not have made it clear, but I started the time at 5PM, that is why I gave bus 1 the extra 80 miles on distance.

OpenStudy (radar):

@loli, any questions?

OpenStudy (radar):

@LivForMusic, would you have worked it differently? I know there are other ways it could be solved.

OpenStudy (anonymous):

@radar thats how i worked it.